OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33

OpenStudy (ivettef365):

first solve the parenthesis by multiplying (3)(x) and (3)(1)

OpenStudy (uri):

Solve the brackets first 3x+3+6=33 3x=33-3-6 3x=24 x=24/3 Divide and answer.